What is the price of Platinum Water Entertainment stock expected to be in 5 years if its annual dividend is expected to grow by 6 percent per year forever, the next dividend is expected in 1 year, the dividend is expected to be 8.63 dollars in 5 years, and the expected return is 11.31 percent per year?
Answer #1

Stock Price in 5 years = Ds(1 + Growth Rate) Rateof Return - Growth Rate

D5 = $ 8.63

Stock Price in 5 years = 8.63(1+0.06) 0.1131 -0.06

Stock Price in 5 years = 9.1478 / 0.0531

Stock Price in 5 years =$ 172.2749

NOTE: The answer to your question has been given below/above. If there is any query regarding the answer, please ask in the comment section. If you find the answer helpful, do upvote. Help us help you.